# Geodesic Distance Between Graphs: A Spectral Metric for Assessing the Stability of Graph Neural Networks

### v1

@@ Abstract
This paper presents a spectral framework for assessing the generalization and stability of Graph Neural Networks (GNNs) by introducing a Graph Geodesic Distance (GGD) metric. For two different graphs with the same number of nodes, our framework leverages a spectral graph matching procedure to find node correspondence so that the geodesic distance between them can be subsequently computed by solving a generalized eigenvalue problem associated with their Laplacian matrices. For graphs with different sizes, a resistance-based spectral graph coarsening scheme is introduced to reduce the size of the bigger graph while preserving the original spectral properties. We show that the proposed GGD metric can effectively quantify dissimilarities between two graphs by encapsulating their differences in key structural (spectral) properties, such as effective resistances between nodes, cuts, the mixing time of random walks, etc. Through extensive experiments comparing with the state-of-the-art metrics, such as the latest Tree-Mover’s Distance (TMD) metric, the proposed GGD metric shows significantly improved performance for stability evaluation of GNNs especially when only partial node features are available.
@@ 1 Introduction
In the era of big data, comparison and distinction between data points are important tasks. A graph is a specific type of data structure that represents the connections between a group of nodes or agents. Comparing two graphs often involves using a pairwise distance measure, where a small distance indicates a high structural similarity and vice versa. To understand the generalization between distribution shifts, it is important to use an appropriate measure of divergence between data distributions, both theoretically and experimentally Chuang et al. [2020]. Determining suitable distance metrics for non-Euclidean data entities, like graphs with or without node attributes, which are fundamental to many graph learning methods such as graph neural networks (GNNs), remains a significant challenge, even though distance metrics for data points in Euclidean space are readily available. The need to develop new analytical techniques that allow the visualization, comparison, and understanding of different graphs has led to a rich field of research study Haslbeck and Waldorp [2018]. This study dives into the exploration of a novel framework for computing geometric distances between graphs, which can be immediately leveraged for many graph-based machine learning (ML) tasks, such as the stability evaluation of GNNs.
@@
Many distance metrics for comparing graphs have previously been proposed Borgwardt et al. [2020]. Some of them are merely based on graph local structures Haussler et al. [1999], Xu et al. [2013], Zhu et al. [2020], Fernández and Valiente [2001], Bunke and Shearer [1998], whereas others exploit both graph structural properties and node attributes Shervashidze et al. [2011], Morris et al. [2019]. For example, the Graph Edit Distance (GED) has been proposed to measure the distance between graphs considering the number of changes needed to match one graph to another Sanfeliu and Fu [1983], Gao et al. [2010], Li et al. [2017]; Distance metrics based on the graph kernel have also been investigated Shervashidze et al. [2011], Vishwanathan et al. [2010], such as the Wasserstein Weisfeiler-Leman metric (WWL) Morris et al. [2019] and the Gromov–Wasserstein metric Mémoli [2011], which allow computing graph distances based on low-dimensional graph representations or optimal transport (OT) Titouan et al. [2019], Chapel et al. [2020], leading to the development of the state-of-the-art graph distance metric called Tree Mover’s Distance (TMD) Chuang and Jegelka [2022].
@@
However, the existing graph distance metrics have distinct limitations. For example, the GED metric can capture local node or edge changes but struggles with global perturbations Sanfeliu and Fu [1983], Gao et al. [2010], Li et al. [2017]; the WWL and TMD metrics heavily rely on node features (attributes) for calculating the distance between graphs, leading to degraded performance when only partial node features are available Rossi et al. [2022], Chen et al. [2022].
@@
To address the limitations of prior methods, we propose a Graph Geodesic Distance (GGD) metric that can be computed through the following steps: for two graphs of the same size (with $n$ nodes), (1) a spectral graph matching procedure is exploited for finding the node correspondence between two graphs, and (2) a geodesic distance metric is defined on the cone of $n\times n$ modified graph Laplacians that can be regarded as a Riemannian manifold Lim et al. [2019]. For graphs of different sizes, we introduce a resistance-based spectral graph coarsening scheme to transform the larger graph into a smaller one suitable for the GGD evaluation. We show that the proposed GGD metric can theoretically capture key structural (spectral) dissimilarities between two graphs, such as mismatches in Laplacian eigenvalues/eigenvectors, cuts, effective-resistance distances, etc.
@@
One distinct advantage of the proposed GGD metric is its capability to compute distances between graphs based on their spectral (structural) properties, while including node feature information into our framework can further improve the accuracy. Therefore, GGD is suitable for analyzing many real-world graphs that may only have partial or even no node features. Moreover, the proposed framework for computing GGDs is more computationally efficient than existing OT-based metrics, such as the TMD metric. Our preliminary results show that the GGD metric has a better correlation with established GNN output compared to the state-of-the-art TMD metric Chuang and Jegelka [2022]. It also allows us to achieve up to $10\%$ accuracy gain and 9×9$\times$9 × runtime speedup in graph classification tasks than TMD.
@@ 2 Existing Graph Distance Metrics
Graph Edit Distance (GED) For non-attributed graph data, a common and simple distance metric is GED. Sanfeliu and Fu [1983], Gao et al. [2010]. Given a set of graph edit operations, also known as elementary graph operations, the GED between two graphs $G_{1}$ and $G_{2}$, written as $\text{GED}(G_{1},G_{2})$, can be defined as:
$\text{GED}(G_{1},G_{2})=\min\limits_{\begin{subarray}{c}(e_{1},\ldots,e_{k})\in\mathcal{P}(G_{1},G_{2})\end{subarray}}\sum_{i=1}^{k}c(e_{i})$ (1)
@@
Tree Mover’s Distance (TMD) TMD is a pseudometric for measuring distances between simple graphs, extending the concept of WWL to multisets of tree structures Chuang and Jegelka [2022]. By progressively adding neighboring nodes to the previous node at each level, we obtain the computation tree of a node. These tree structures are crucial in graph analysis Weisfeiler and Leman [1968], Pearson [1905] and graph kernels Ramon and Gärtner [2003], Shervashidze et al. [2011]. TMD uses hierarchical optimal transport (HOT) to analyze these computational trees from input graphs. For a graph $G=(V,E)$ with node features $f_{v}\in\mathbb{R}^{s}$ for node $v\in V$, let $T_{v}^{1}=v$, and $T_{v}^{L}$ be the depth-L computation tree of node $v$. The multiset of these trees for $G$ is $T_{G}^{L}=\{T_{v}^{L}\}_{v\in V}$. The number and shape of trees must match to calculate optimal transport between two multisets of trees. If multisets are uneven, they are augmented with blank nodes. For multisets $T_{p}$ and $T_{q}$, the augmenting function $\sigma$ adds blank trees to equalize their sizes. A blank tree $T_{\mathbb{O}}$ has a single node with a zero vector feature $\mathbb{O}_{p}\in\mathbb{R}^{s}$:
$\sigma:(T_{p},T_{q})\rightarrow\left(T_{p}\cup T_{\mathbb{O}}^{\max(|T_{q}|-|T_{p}|,0)},T_{q}\cup T_{\mathbb{O}}^{\max(|T_{p}|-|T_{q}|,0)}\right)$ (2)
@@
$\mathrm{OT}_{d}(X,Y):=\min_{\gamma\in\Gamma(X,Y)}\langle C,\gamma\rangle\quad\Gamma(X,Y)=\left\{\gamma\in\mathbb{R}_{+}^{m\times m}\mid\gamma\mathbb{1}_{m}=\gamma^{\top}\mathbb{1}_{m}=\mathbb{1}_{m}\right\}$ (3)
Here $\Gamma$ is the set of transportation plans that satisfies the flow constrain $\gamma\mathbb{1}_{m}=\gamma^{\top}\mathbb{1}_{m}=\mathbb{1}_{m}$ Chuang and Jegelka [2022].
@@
$\text{TD}_{w}\left(T_{p},T_{q}\right):=\begin{cases}\left\|f_{r_{p}}-f_{r_{q}}\right\|+w(L)\cdot\mathrm{OT}_{\mathrm{TD}_{w}}\left(\sigma\left(T_{r_{p}},T_{r_{q}}\right)\right)&\text{if }L>1\\ \left\|f_{r_{p}}-f_{r_{q}}\right\|&\text{otherwise}\end{cases}$ (4)
@@
$\text{TMD}_{w}^{L}(G_{1},G_{2})=\text{OT}_{\mathrm{TD}_{w}}(\sigma(\mathbf{T}_{G_{1}}^{L},\mathbf{T}_{G_{2}}^{L}))$ (5)
@@ 3 GGD: A Geodesic Distance Metric for Graphs
Laplacian matrices on the Riemannian manifold
To overcome the limitations of the previous OT-based graph distance metrics Chuang and Jegelka [2022], we aim to develop a new pseudometric that can effectively encapsulate topological differences between graphs. One way to represent a simple graph is through its Laplacian matrix, which is a Symmetric Positive Semidefinite matrix. Adding a small positive value to each diagonal element will allow us to transform the original Laplacian matrix into a Symmetric Positive Definite (SPD) one, which is called the modified Laplacian matrix in this work. Then we can consider the cone of such modified Laplacian matrices as a natural Riemannian manifold Lim et al. [2019], where each modified Laplacian has the same dimensions (same number of rows/columns) and can be regarded as a data point on the Riemannian manifold Vemulapalli and Jacobs [2015], Pennec et al. [2006]. In the last, the geodesic distance between two graphs can be defined as the shortest path distance on the Riemannian manifold if their node correspondence is known in advance. This approach is more appropriate than directly comparing the graphs in the Euclidean space Lim et al. [2019], Crane et al. [2020], Huang et al. [2015]. We will later show (Section 4.3) that such a geodesic distance metric can effectively capture structural (spectral) mismatches between graphs.
A spectral framework for computing GGDs
Before computing GGDs, it is necessary to establish the node-to-node correspondence between two graphs. This can be achieved by leveraging existing graph-matching techniques Livi and Rizzi [2013], Emmert-Streib et al. [2016], Caetano et al. [2009]. In this work, we will leverage a recent spectral graph matching framework that has been shown to recover accurate matching with high probability Fan et al. [2020].
The proposed GGD metric can be computed in the following two phases. Phase 1 consists of a spectral graph matching step, using combinatorial optimization with the eigenvalues/eigenvectors of the graph adjacency matrices to identify the approximate node-to-node correspondence. Phase 2 computes the GGD between the modified Laplacian matrices of the matched graphs by exploiting generalized eigenvalues. The proposed GGD metric differs from previous OT-based graph distance metrics in its ability to accurately represent structural discrepancies between graphs, enabling us to uncover the topological variations between them more effectively. Since only the graph Laplacian (adjacency) matrix is required to calculate the GGD, our metric can even work effectively for graphs without node feature information.
@@
Table 1: Distance between graphs with simple perturbation
@@
row1 = c, row2 = c, cell11 = r=2, cell12 = c=4, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, hline1 = -0.08em, hline2 = 2-50.03em, hline3 = -, hline6 = -0.08em, Graph
Pairs & Distance Metrics (Normalized)
GGD TMD with
NF, L = 4 TMD without
NF, L = 4
GED
@@ 4.1 Phase 1: Spectral Graph Matching for Finding Node Correspondence
Computing the GGD metric between two input graphs requires solving a graph matching problem in advance. Graph matching techniques can be used to establish node-to-node correspondence by seeking a bijection between node sets to maximize the alignment of edge sets Livi and Rizzi [2013], Emmert-Streib et al. [2016], Caetano et al. [2009]. This combinatorial optimization problem can be cast into a Quadratic Assignment Problem, which is NP-hard to solve or approximate Fan et al. [2020], Wang et al. [2020].
Figure 1: Graphs with simple perturbations
In this study, we will exploit a spectral graph matching method called GRAMPA (GRAph Matching by Pairwise eigen-Alignments) Fan et al. [2020] to find the approximate node correspondence between two graphs. GRAMPA starts with comparing the eigenvectors of the adjacency matrices of the input graphs. Instead of comparing only the eigenvectors responding to the largest eigenvalues, it considers all pairs of eigenvectors/eigenvalues to generate a similarity matrix. This similarity matrix can be constructed by summing up the outer products of eigenvector pairs, weighted by a Cauchy kernel Fan et al. [2020]. Subsequently, a rounding procedure will be performed to determine the optimal match between nodes employing the similarity matrix.
@@
$\widehat{X}=\sum_{i,j=1}^{n}w\left(\zeta_{i},\mu_{j}\right)\cdot u_{i}u_{i}^{\top}\mathbf{J}v_{j}v_{j}^{\top},\quad\text{where}\quad w(x,y)=\frac{1}{(x-y)^{2}+\eta^{2}}$ (6)
@@
$\widehat{\pi}={\operatorname{argmax}}\sum_{i=1}^{n}\widehat{X}_{i,\pi(i)}$ (7)
This LAP can be solved efficiently using the Hungarian Algorithm Fan et al. [2020], Kuhn [1955], where:
@@
Given symmetric matrices $A$, $B$ and $Z$ from the Gaussian Wigner model, where $B_{\pi^{*}}=A+\sigma Z$, there exist constants $c,c^{\prime}>0$ such that if $1/n^{0.1}\leq\eta\leq c/\log n$ and $\sigma\leq c^{\prime}\eta$, then with probability at least $1-n^{-4}$, GRAMPA Algorithm correctly recovers the permutation matrix $\pi^{*}$ from the Similarity matrix $\widehat{X}$ Fan et al. [2020]. Its proof can be found in the supporting documents A.2.
@@
$\textnormal{Best Match to }B=\widehat{\pi}\cdot A\cdot\widehat{\pi}^{\top},\quad\textnormal{Best Match to }A=\widehat{\pi}^{\top}\cdot B\cdot\widehat{\pi}$ (9)
In practice, the graph matching performance is not too sensitive to the choice of tuning parameter $\eta$. For small-sized graphs, such as the MUTAG datasetMorris et al. [2020], setting $\eta=0.5$ yields satisfactory results in matching. In Appendix A.3, the effect of $\eta$ for computing GGDs has been comprehensively analyzed.

### v2

@@ Abstract
This paper introduces a spectral framework for assessing the generalization of Graph Neural Networks (GNNs) through a new Graph Geodesic Distance (GGD) metric. For two different graphs with the same number of nodes, our framework leverages a spectral graph matching procedure to find node correspondence so that the geodesic distance between them can be subsequently computed by solving a generalized eigenvalue problem associated with their Laplacian matrices. For graphs of different sizes, a resistance-based spectral graph coarsening scheme is introduced to reduce the size of the larger graph while preserving the original spectral properties. We show that the proposed GGD metric can effectively quantify dissimilarities between two graphs by encapsulating their differences in key structural (spectral) properties, such as effective resistances between nodes, cuts, the mixing time of random walks, etc. Through extensive experiments comparing with the state-of-the-art metrics, such as the latest Tree-Mover’s Distance (TMD) metric, the proposed GGD metric shows significantly improved performance for graph classification and stability evaluation of GNNs, especially when only partial node features are available.
@@ 1 Introduction
In the era of big data, comparison and distinction between data points are important tasks. A graph is a specific type of data structure that represents the connections between a group of nodes or agents. Comparing two graphs often involves using a pairwise distance measure, where a small distance indicates a high structural similarity and vice versa. To understand the generalization between distribution shifts, it is important to use an appropriate measure of divergence between data distributions, both theoretically and experimentally (Chuang et al., 2020). Determining suitable distance metrics for non-Euclidean data entities, like graphs with or without node attributes, which are fundamental to many graph learning methods such as graph neural networks (GNNs), remains a significant challenge, even though distance metrics for data points in Euclidean space are readily available. The need to develop new analytical techniques that allow the visualization, comparison, and understanding of different graphs has led to a rich field of research study (Haslbeck & Waldorp, 2018). This study dives into the exploration of a novel framework for computing geometric distances between graphs, which can be immediately leveraged for many graph-based machine learning (ML) tasks, such as graph classification or the stability evaluation of GNNs.
@@
Many distance metrics for comparing graphs have previously been proposed (Borgwardt et al., 2020). Some of them are merely based on graph local structures (Tam & Dunson, 2022; Haussler et al., 1999; Xu et al., 2013; Zhu et al., 2020; Fernández & Valiente, 2001; Bunke & Shearer, 1998), whereas others exploit both graph structural properties and node attributes (Shervashidze et al., 2011; Morris et al., 2019). For example, the Graph Edit Distance (GED) has been proposed to measure the distance between graphs considering the number of changes needed to match one graph to another (Sanfeliu & Fu, 1983; Gao et al., 2010; Li et al., 2017); Distance metrics based on the graph kernel have also been investigated (Shervashidze et al., 2011; Vishwanathan et al., 2010), such as the Wasserstein Weisfeiler-Leman metric (WWL) (Morris et al., 2019) and the Gromov–Wasserstein metric (Mémoli, 2011), which allow computing graph distances based on low-dimensional graph representations or optimal transport (OT) (Titouan et al., 2019; Chapel et al., 2020), leading to the development of the state-of-the-art graph distance metric called TMD (Chuang & Jegelka, 2022).
@@
However, the existing graph distance metrics have distinct limitations. For example, the GED metric can capture local node or edge changes but struggles with global perturbations (Sanfeliu & Fu, 1983; Gao et al., 2010; Li et al., 2017); the WWL and TMD metrics heavily rely on node features (attributes) for calculating the distance between graphs, leading to degraded performance when only partial node features are available (Rossi et al., 2022; Chen et al., 2022).
@@
To address the limitations of prior methods, we propose a Graph Geodesic Distance (GGD) metric that can be computed through the following steps: for two graphs of the same size (with $n$ nodes), (1) a spectral graph matching procedure is exploited for finding the node correspondence between two graphs, and (2) a geodesic distance metric is defined on the cone of $n\times n$ modified graph Laplacians that can be regarded as a Riemannian manifold (Lim et al., 2019). For graphs of different sizes, we introduce a resistance-based spectral graph coarsening scheme to transform the larger graph into a smaller one suitable for the GGD evaluation. We show that the proposed GGD metric can theoretically capture key structural (spectral) dissimilarities between two graphs, such as mismatches in Laplacian eigenvalues/eigenvectors, cuts, effective-resistance distances, etc.
@@
One distinct advantage of the proposed GGD metric is its capability to compute distances between graphs based on their spectral (structural) properties, while including node feature information into our framework can further improve the accuracy. Therefore, GGD is suitable for analyzing many real-world graphs that may only have partial or even no node features. Moreover, the proposed framework for computing GGDs is more computationally efficient than existing OT-based metrics, such as the TMD metric.
Our empirical results show that GGD can effectively measure the dissimilarities between graphs: (1) support vector classifiers (SVC) using GGDs perform competitively with state-of-the-art GNN models and graph kernels on graph classification benchmarks; (2) we demonstrate that the GGD metric allows us to quantify the stability of GNN models for graph classification tasks by checking whether two graphs with a small GGD will lead to a significant dissimilarity in the GNN output embeddings. We also show that the GGD metric has a better correlation with established GNN outputs compared to the state-of-the-art TMD metric (Chuang & Jegelka, 2022) when only partial node features are available: up to a $10\%$ accuracy gain and a 9×9$\times$9 × runtime speedup have been achieved in various graph classification tasks.
@@ 2 Existing Graph Distance Metrics
Graph Edit Distance (GED) For non-attributed graph data, a common and simple distance metric is GED. (Sanfeliu & Fu, 1983; Gao et al., 2010). Given a set of graph edit operations, also known as elementary graph operations, the GED between two graphs $G_{1}$ and $G_{2}$, written as $\text{GED}(G_{1},G_{2})$, can be defined as:
$\text{GED}(G_{1},G_{2})=\min\limits_{\begin{subarray}{c}(e_{1},\ldots,e_{k})\in\mathcal{P}(G_{1},G_{2})\end{subarray}}\sum_{i=1}^{k}c(e_{i}),$ (1)
@@
Tree Mover’s Distance (TMD) TMD is a pseudometric for measuring distances between simple graphs, extending the concept of WWL to multisets of tree structures (Chuang & Jegelka, 2022). By progressively adding neighboring nodes to the previous node at each level, we obtain the computation tree of a node. These tree structures are crucial in graph analysis (Weisfeiler & Leman, 1968; Pearson, 1905) and graph kernels (Ramon & Gärtner, 2003; Shervashidze et al., 2011). TMD uses hierarchical optimal transport (HOT) to analyze these computational trees from input graphs. For a graph $G=(V,E)$ with node features $f_{v}\in\mathbb{R}^{s}$ for node $v\in V$, let $T_{v}^{1}=v$, and $T_{v}^{L}$ be the depth-L computation tree of node $v$. The multiset of these trees for $G$ is $T_{G}^{L}=\{T_{v}^{L}\}_{v\in V}$. The number and shape of trees must match to calculate optimal transport between two multisets of trees. If multisets are uneven, they are augmented with blank nodes. For multisets $T_{p}$ and $T_{q}$, the augmenting function $\sigma$ adds blank trees to equalize their sizes. A blank tree $T_{\mathbb{O}}$ has a single node with a zero vector feature $\mathbb{O}_{p}\in\mathbb{R}^{s}$:
$\sigma:(T_{p},T_{q})\rightarrow\left(T_{p}\cup T_{\mathbb{O}}^{\max(|T_{q}|-|T_{p}|,0)},T_{q}\cup T_{\mathbb{O}}^{\max(|T_{p}|-|T_{q}|,0)}\right).$ (2)
@@
$\mathrm{OT}_{d}(X,Y):=\min_{\gamma\in\Gamma(X,Y)}\langle C,\gamma\rangle,\quad\Gamma(X,Y)=\left\{\gamma\in\mathbb{R}_{+}^{m\times m}\mid\gamma\mathbb{1}_{m}=\gamma^{\top}\mathbb{1}_{m}=\mathbb{1}_{m}\right\}.$ (3)
Here $\Gamma$ is the set of transportation plans that satisfies the flow constrain $\gamma\mathbb{1}_{m}=\gamma^{\top}\mathbb{1}_{m}=\mathbb{1}_{m}$. (Chuang & Jegelka, 2022).
@@
$\text{TD}_{w}\left(T_{p},T_{q}\right):=\begin{cases}\left\|f_{r_{p}}-f_{r_{q}}\right\|+w(L)\cdot\mathrm{OT}_{\mathrm{TD}_{w}}\left(\sigma\left(T_{r_{p}},T_{r_{q}}\right)\right),&\text{if }L>1\\ \left\|f_{r_{p}}-f_{r_{q}}\right\|,&\text{otherwise}\end{cases}$ (4)
@@
$\text{TMD}_{w}^{L}(G_{1},G_{2})=\text{OT}_{\mathrm{TD}_{w}}(\sigma(\mathbf{T}_{G_{1}}^{L},\mathbf{T}_{G_{2}}^{L})).$ (5)
@@ 3 GGD: A Geodesic Distance Metric for Graphs
Modified Laplacian matrices on the Riemannian manifold
To overcome the limitations of previous OT-based graph distance metrics (Chuang & Jegelka, 2022), we aim to develop a new metric that effectively captures topological differences between graphs. One way to represent a simple graph is through its Laplacian matrix, which is a Symmetric Positive Semidefinite (SPSD) matrix. Graph representation using adjacency and Laplacian matrices is briefly discussed in Appendix A.1. Adding a small positive value to each diagonal element allows us to transform the original Laplacian matrix into a Symmetric Positive Definite (SPD) matrix, which is referred to as the Modified Laplacian Matrix in this work. In Appendix A.6, we describe the effect of this small value on the GGD calculation. We can then consider the cone of such modified Laplacian matrices as a natural Riemannian manifold (Lim et al., 2019), where each modified Laplacian, having the same dimensions (same number of rows/columns), can be regarded as a data point on this Riemannian manifold (Vemulapalli & Jacobs, 2015; Pennec et al., 2006). Details about the Riemannian manifold are provided in Appendix A.2. Finally, the geodesic distance between two graphs can be defined as the shortest path distance on the Riemannian manifold, assuming their node correspondence is known in advance. This approach is more appropriate than directly comparing the graphs in Euclidean space (Lim et al., 2019; Crane et al., 2020; Huang et al., 2015). We will later demonstrate (Section 4.3) that such a geodesic distance metric can effectively capture structural (spectral) mismatches between graphs.
A two-phase spectral framework for computing GGDs
Before computing GGDs, it is necessary to establish the node-to-node correspondence between two graphs. This can be achieved by leveraging existing graph-matching techniques (Livi & Rizzi, 2013; Emmert-Streib et al., 2016; Caetano et al., 2009). In this work, we will leverage a recent spectral graph matching framework that has been shown to recover accurate matching with high probability (Fan et al., 2020). The proposed GGD metric can be computed in the following two phases. Phase 1 consists of a spectral graph matching step, using combinatorial optimization with the eigenvalues/eigenvectors of the graph adjacency matrices to identify the approximate node-to-node correspondence. Phase 2 computes the GGD between the modified Laplacian matrices of the matched graphs by exploiting generalized eigenvalues. The proposed GGD metric differs from previous OT-based graph distance metrics in its ability to accurately represent structural discrepancies between graphs, enabling us to uncover the topological variations between them more effectively. Since only the graph Laplacian (adjacency) matrix is required to calculate the GGD, our metric can even work effectively for graphs without node feature information.
@@
Table 1: Distance between graphs with simple perturbations.
@@
roweven = c, row3 = c, row5 = c, cell11 = r=2, cell12 = c=4c, hline1,3,6 = -, hline2 = 2-5, Graph pairs & Distance metrics (Normalized)
GGD TMD with NF, L = 4 TMD without NF, L = 4 GED
@@ 4.1 Phase 1: Spectral Graph Matching for Finding Node Correspondence
Computing the GGD metric between two input graphs requires solving a graph-matching problem in advance. Without knowing the node-to-node correspondence that can be achieved through a graph-matching step, the distance between modified Laplacian matrices may be significantly higher than the minimum possible distance. In this work, we aim to find the infimum between two SPD matrices on the Riemannian manifold formed by modified Laplacian matrices, which can be accomplished through a graph-matching phase. Graph matching techniques can be used to establish node-to-node correspondence by seeking a bijection between node sets to maximize the alignment of edge sets (Livi & Rizzi, 2013; Emmert-Streib et al., 2016; Caetano et al., 2009). This combinatorial optimization problem can be cast into a Quadratic Assignment Problem, which is NP-hard to solve or approximate (Fan et al., 2020; Wang et al., 2020).
Figure 1: Graphs with simple perturbations.
In this study, we will exploit a spectral graph matching method called GRAMPA (GRAph Matching by Pairwise eigen-Alignments) (Fan et al., 2020) to find the approximate node correspondence between two graphs. GRAMPA starts with comparing the eigenvectors of the adjacency matrices of the input graphs. Instead of comparing only the eigenvectors responding to the largest eigenvalues, it considers all pairs of eigenvectors/eigenvalues to generate a similarity matrix. This similarity matrix can be constructed by summing up the outer products of eigenvector pairs, weighted by a Cauchy kernel (Fan et al., 2020). Subsequently, a rounding procedure will be performed to determine the optimal match between nodes employing the similarity matrix.
@@
$\widehat{X}=\sum_{i,j=1}^{n}w\left(\zeta_{i},\mu_{j}\right)\cdot u_{i}u_{i}^{\top}\mathbf{J}v_{j}v_{j}^{\top},\text{where }w(x,y)=\frac{1}{(x-y)^{2}+\eta^{2}}.$ (6)
@@
$\widehat{\pi}={\operatorname{argmax}}\sum_{i=1}^{n}\widehat{X}_{i,\pi(i)},$ (7)
which can be efficiently solved using the Hungarian algorithm (Fan et al., 2020). However, one simpler rounding procedure was advocated in (Fan et al., 2020) with theoretical results supporting the rounding procedure, which is given by the following equation:
@@
here the permutation estimate matrix is constructed by selecting the largest index from each row. While LAP provides optimal matching, its computational complexity can become expensive for very large graphs. By carefully choosing $\eta$, the same match recovery holds if rounding is performed using equation 7 instead of solving the LAP in equation 8 (Fan et al., 2020).
@@
Given symmetric matrices $A_{1}$, $A_{2}$ and $Z$ from the Gaussian Wigner model, where $A_{2\pi^{*}}=A_{1}+\sigma Z$, there exist constants $c,c^{\prime}>0$ such that if $1/n^{0.1}\leq\eta\leq c/\log n$ and $\sigma\leq c^{\prime}\eta$, then with probability at least $1-n^{-4}$, GRAMPA Algorithm correctly recovers the permutation matrix $\pi^{*}$ from the Similarity matrix $\widehat{X}$ (Fan et al., 2020). Its proof can be found in the supporting documents A.5.
@@
$\textnormal{Best Match to }A_{2}=\widehat{\pi}A_{1}\widehat{\pi}^{\top},\quad\textnormal{Best Match to }A_{1}=\widehat{\pi}^{\top}A_{2}\widehat{\pi}.$ (9)
In practice, the graph matching performance is not too sensitive to the choice of tuning parameter $\eta$. For small-sized graphs, such as the MUTAG dataset(Morris et al., 2020), setting $\eta=0.5$ yields satisfactory results in matching. In Appendix A.9, the effect of $\eta$ for computing GGDs has been comprehensively analyzed.